OpenStudy (anonymous):

Verify the trig identity: 2cos^2(x/2)=(sin^2x)/(1-cosx)

OpenStudy (slaaibak):

Start with lefthandside: 2cos^2(x/2) = 1 + cos x = (1 + cosx)(1-cosx)/(1-cosx) = (1 - cos^2x)/(1+cosx) = (sin^2 x)/(1-cosx)

OpenStudy (anonymous):

thank you so much!

OpenStudy (slaaibak):

:)